OpenStudy (anonymous):

Each side of a triangle is less than 1 ft long. Can its area be greater than 100 sq feet?

OpenStudy (cwrw238):

what do you think?

OpenStudy (anonymous):

I think its not possible but am not able to give a specific reason

OpenStudy (cwrw238):



OpenStudy (cwrw238):

the above triangle has all sides equal to 1 and is bigger in area than any triangle with each side less than 1 its area is 1 * 0.5 = 0.5 square feet that proves it

OpenStudy (whpalmer4):

Formula for the area of a triangle is \[A = \frac{1}{2}bh\]Even if we assume the height could be \(1 \text{ ft}\), which it clearly cannot be, the maximum area we could get would be \[A = \frac{1}{2}(1\text{ ft})(1\text{ ft}) = \frac{1}{2}\text{ ft}^2\]and that is much less than \(100 \text{ ft}^2\)!

OpenStudy (mathstudent55):



OpenStudy (anonymous):

Thank you all for all your help! You have all all been so supportive.....God bless u all.....thanx again!